### v2

# Demonstrating dynamic surface codes

Abstract
A remarkable characteristic of quantum computing is the potential for reliable computation despite faulty qubits. This can be achieved through quantum error correction, which is typically implemented by repeatedly applying static syndrome checks, permitting correction of logical information [1, 2, 3]. Recently, the development of time-dynamic approaches to error correction has uncovered new codes [4] and new code implementations [5, 6, 7]. In this work, we experimentally demonstrate three time-dynamic implementations of the surface code, each offering a unique solution to hardware design challenges and introducing flexibility in surface code realization. First, we embed the surface code on a hexagonal lattice, reducing the necessary couplings per qubit from four to three. Second, we walk a surface code, swapping the role of data and measure qubits each round, achieving error correction with built-in removal of accumulated non-computational errors. Finally, we realize the surface code using iSWAP gates instead of the traditional CNOT, extending the set of viable gates for error correction without additional overhead. We measure the error suppression factor when scaling from distance-3 to distance-5 codes of , , and , achieving state-of-the-art error suppression for each. With detailed error budgeting, we explore their performance trade-offs and implications for hardware design. This work demonstrates that dynamic circuit approaches satisfy the demands for fault-tolerance and opens new alternative avenues for scalable hardware design [8].
Quantum Error Correction (QEC) enables accurate quantum computation in the presence of uncontrolled physical noise [9]. The surface code [1, 10, 11] is a preeminent example, capable of suppressing physical errors below a relatively high threshold and providing excellent capabilities for logical computation [2, 12]. These desirable characteristics have inspired extensive experimental work on the surface code [13, 14], including recently demonstrated below-threshold performance on a superconducting quantum processor [15] and transversal logical gates in a neutral atom quantum processor [16]. To date, all of these experiments have been designed around a square grid of statically assigned qubits implementing CNOT/CZ entangling interactions to match the standard circuit implementation. This was assumed to comprise the least demanding experimental implementation of the code, since fault-tolerant circuits are traditionally constructed from static QEC codes.
Recent progress in the theory of time-dynamic QEC, including dynamic codes [4] and space-time stabilizers [17, 18], challenge the traditional static stabilizer approach and unlock new opportunities for the construction of fault-tolerant circuits. The new framework of the space-time detecting region generalizes these approaches, defined by local and connected regions of space-time that report the presence of errors. This framework permits the re-interpretation of QEC circuits as intermeshed detecting regions, featuring overlapping sensitivity such that all physical errors can be appropriately triangulated. It also highlights that maintaining the precise shape of such detecting regions is less crucial than ensuring all circuit locations are covered. Exploiting this freedom opens the door to an expansive equivalence class of circuits for implementing well-understood QEC codes [5, 19, 6, 20, 7], providing distinct flexibility in their hardware implementation.
In this work, we demonstrate the ability to switch out the circuit implementation of the surface code, achieving error suppression from distance-3 to distance-5 with three time-dynamic circuits that feature qualitatively different demands on the underlying hardware. These circuits are examples representing particularly large departures from previous assumptions about the hardware and gateset necessary to implement the surface code. First, we embed the surface code onto a hexagonal grid, reducing the necessary qubit connectivity from four neighbours to three. Second, we perform a walking circuit in which the qubits are dynamically reassigned between data and measure, frustrating the spread of time-correlated errors. Finally, we perform the surface code using iSWAP gates rather than CNOT/CZ gates, demonstrating that QEC can be performed using entangling interactions previously reserved for NISQ experiments. We operate all three implementations on Willow processors [15], optimised for the standard surface code circuit, permitting direct comparisons of performance but leaving open the possibility of optimising future hardware around alternative circuits.
The newfound freedom in circuit implementation can be understood naturally using detecting regions, which indicate all locations in a QEC circuit where a non-commuting error would be detected by measurements the region touches. Fig. 1a shows a standard distance-5 surface code patch, where we highlight a single detecting region in Fig. 1c. This Z-type detecting region is sensitive to any X or Y error occurring in highlighted locations in the circuit. The region propagates through the circuit in the same way that Pauli terms propagate through Clifford operations, illustrated in Fig. 1b, with more details given in [5]. The detecting region persists over two consecutive cycles, first expanding from a single qubit to cover the static code stabilizer immediately after the first measurement gate, which we call the end-cycle state, then contracting to end in a second measurement. Time-slices of the detecting region are shown in Fig. 1d; using a tiling of spacetime through overlapping detecting regions, we ensure that all relevant circuit locations are checked for errors. Given this general framework, circuit flexibility arises by modifying the gate layers and associated detecting regions while preserving the tiling of spacetime.
The end-cycle states for the three alternative constructions measured in this work are illustrated in Fig. 1e. Each alternative construction features unique alternating end-cycle states, as opposed to the standard code which has a static end-cycle state. Such periodic deformation of the detecting region tiling allows the alternative circuit constructions. In this manuscript, we measure the QEC performance of these three time-dynamic circuits in isolation using a logical memory experiment, proving their ability to reduce the logical error rate as the surface code distance is increased from 3 to 5, while discussing their unique hardware requirements.
Hexagonal lattice Surface Code
The assumed necessity of a square lattice for the surface code has been a driving influence on superconducting device layouts, with experiments to date implementing four local couplings on each qubit [13, 21, 14, 15]. This strong device requirement adds substantial design complexity and has limited some architectures from implementing the surface code [22, 23, 24].
Using a time-dynamic circuit featuring an alternative detecting region tiling, a recent proposal provides an embedding of the surface code on a hexagonal grid [5]. In Fig. 2a we highlight the key insights of this proposal. The first insight is that the last two-qubit gate layer of each QEC cycle can be altered to match the first layer with the CNOT orientation inverted, using the same qubit-qubit coupling layer twice in each cycle. This results in the expanding detecting region being shifted laterally while preserving the square weight-4 pattern of stabilizers. The second key insight is to use a time-reversed circuit every second cycle, refocusing the detecting region to its starting position during the contraction by an opposite lateral shift. This pattern of two opposing detecting regions is sufficient to tile the space-time bulk of the circuit. We include details on the construction of detecting regions at the spatial boundaries in (see SM sec. C1).
We benchmark the hexagonal implementation using memory experiments run on a superconducting quantum processor. We first prepare the or logical state, apply a variable number of error correction cycles, and finally measure the logical qubit in the initialized logical basis. We use the same 72-qubit device as in Ref. [15], which features flux-tunable capacitive couplers [25, 26]. We tune the unused couplers to their zero-coupling bias. The distance-5 hexagonal lattice circuit uses 49 computational qubits and 64 couplers, 20% fewer than the 80 couplers used in the standard implementation. Even with fewer physical couplers used, the bulk detecting regions of both standard and hexagonal lattice implementations touch 22 CNOT gates, leading to similar sensitivity to two-qubit gate errors (see SM sec. B). We optimize the operating frequencies of the gates taking into account the geometry of the circuit as in Ref. [27]. Additionally, we use the Data Qubit Leakage Removal (DQLR) technique developed in Ref. [28] to mitigate the impact of leakage on the logical performance.
We run the memory experiment for a distance-5 hexagonal surface code and five tiling distance-3 codes. In Fig. 2b we report the logical error probability at varying circuit depths, along with the logical error rate per cycle for different decoders (Fig. 2c), ordered from fastest to most accurate. Using the Harmony decoder [29] with a noise model prior trained using reinforcement learning [30], we measure a distance-5 logical error rate of and a mean distance-3 logical error rate of (averaged over five codes), fit with the same procedure described in [15]. From this, we extract an error suppression factor of , indicating the surface code can be operated on a hexagonal lattice with performance matching the standard square lattice implementation () [15].
To build confidence in our results, we measure the fidelity of each component of the error correction circuit through independent benchmarking. From these benchmarks and the circuit’s detecting region description, we analytically compute a linear error budget for each region’s detection probability, shown in Fig. 2d and accurate to first order in physical error probabilities (see SM sec. A4). We validate this budget by computing the root-mean-square (RMS) between the analytic prediction and measured value for detection probabilities, finding a RMS of for the distance-5 code over 54 cycles (see SM sec. C3). In Fig. 2e, we summarize the distribution of different physical error contributions to the detection probabilities. The dominant source of errors are the entangling gate (CZ), the idling error on data qubits during measurement and reset (Dynamical Decoupling, or DD), and readout (RO). Since the boundary detectors use twice-fewer entangling gates, the CZ contribution is halved. Single-qubit gates, leakage removal, and reset combined make up only 15% of the detection budget, highlighting the quality of these operations.
Walking Surface Code
In the standard circuit construction for a QEC code, a static set of data qubits are used to support logical information, and static measure qubits are added to perform the parity checks. The strict distinction between these two categories has a direct consequence in hardware implementations. For example, the data qubits always hold important logical information and therefore are never replaced or reset to their ground state , allowing them to accumulate non-computational errors. In superconducting circuits, leakage to and higher represents one such accumulating error, limiting experimental performance [31, 13] and requiring hardware intervention [32, 28, 33, 34]. In cold atom implementations, qubit loss when the atoms escape the trap is another critical example of an accumulating error [16, 35, 36, 37]. In the framework of detecting regions demonstrated here, the distinction between measure and data qubits is not necessary. It has been proposed that the surface code can be implemented using a circuit where each qubit alternates each cycle between the measure and data roles [5, 20]. These circuits also shift the physical qubit support of the logical qubit and permit movement through space, so we refer to them as walking circuits.
In a walking circuit, each physical qubit experiences a measurement and an opportunity for reset or replacement every two cycles. For superconducting qubits, if a multi-level reset strategy is used at each cycle, the walking circuit presents the advantage of periodically removing leakage from all qubits. For neutral atoms, it relaxes the necessary atom lifetime between opportunities for loss detection and replacement, limiting the errors induced by a loss event and allowing rapid atom replacement with a minimal impact on logical performance.
Walking is permitted by the gate layers and detecting region time-slices shown in Fig. 3a. Using this walking circuit, we benchmark the scaling of a logical memory on a 105-qubit superconducting quantum processor [15] (a different device than was used for the hexagonal implementation), with logical error probability results shown in Fig. 3b and logical error rate per cycle for different decoders in Fig. 3c. Here, we realize the distance-5 and embedded distance-3 walking circuits without any explicit leakage removal strategy other than multi-level reset (no DQLR). In order to allow room for walking, the distance-5 circuit uses 58 qubits, more than the standard circuit’s 49, leading to a slight overhead in logical error rate. For a logical memory experiment, we measure distance-5 and average distanced-3 logical error rates of , , leading to an error suppression factor . The median CZ Pauli error of used in the walking circuit is larger than the hexagonal implementation (see SM sec. D2), leading to a reduction in .
To benchmark the walking circuit’s leakage dynamics, we measure the leakage population at each time-slice of the circuit, separated into the two sub-grids of qubits as shown in Fig. 3d. At each moment when a reset is applied to a group of qubits, the leakage population on that group drops by around an order in magnitude, followed by a ramp in population until they are reset again two cycles later. The staggered multi-level reset applications limit leakage population buildup over many cycles, preventing a rising contribution to the overall logical performance as the code continues.
Leakage also causes time-correlated errors that are not properly handled by the decoder. We directly probe these correlations by measuring the detector autocorrelation in Fig. 3e, which is the average covariance of the detecting events separated by cycles. While two-round correlations are expected from time-like Pauli errors (e.g. readout or reset errors), this measurement reveals longer time correlations induced by non-Pauli errors. Without leakage removal in the standard code, we measure a long correlation tail lasting nearly 50 cycles, while the walking circuit and the standard code with leakage removal significantly reduce these long-time correlations.
In addition to these leakage measurements, we apply the same detector budgeting techniques introduced in the previous section to reveal the contributions of each physical error mechanism on detection probabilities in the circuit (Fig. 3f). Our Pauli model produces detections that have a weaker correlation with experiment (RMS , see (see SM sec. D3)) than the hexagonal implementation, indicating that some error mechanisms are not fully captured by the Pauli model, but reveals the same dominant contributions as the hexagonal implementation. The CZ contribution is larger as expected from the reduced gate performance. Even with the reduction in logical performance and , our results demonstrate that a walking surface code can suppress errors as distance increases with a built-in leakage removal, paving the way for new architectures that can use this property to their advantage.
iSWAP Surface Code
The third implementation explored in this work uses the iSWAP family of gates for the entangling layers instead of the standard CNOT/CZ family. It was previously unclear how to implement quantum error correction using the iSWAP gate, as it distorts the stabilizers, preventing them from returning to their initial static code state. However, using the dynamic circuit technique of time-reversing every second cycle, we can find complementary distortions of the stabilizers that permit the iSWAP gate to preserve the structure of the surface code. In Fig. 4a, we show the expansion and contraction of a single detecting region. The end-cycles forms a distorted arrowhead shape instead of the usual square. Nonetheless, the arrowhead pattern can be clearly identified as a two-colorable checkerboard of weight-4 X and Z stabilizers, sufficient for QEC of the surface code.
In Fig. 4b and Fig. 4c, we benchmark the logical error rate of a distance-5 and embedded distance-3 iSWAP implementation of surface codes on our 72-qubit device (the distance-5 circuit uses 57 qubits). Using decoder H*, we measure logical error rates of and averaged over the five distance 3 codes, here again without using DQLR (relying on multi-level reset as the only active source of leakage removal). We extract the error suppression factor , demonstrating the viability of a surface code implemented with iSWAP gates.
In previous experiments, the so-called Sycamore gate using a short, on-resonance interaction has been used as an entangling resource to realize beyond-classical circuits [38, 39]. The Sycamore gate can be decomposed as an iSWAP gate and a controlled-phase gate, with arbitrary angle (see SM sec. E2). While this controlled-phase is not a problem for NISQ experiments [40], it adds a Pauli error per gate of for stabilizer codes that require a strict Clifford gate, as confirmed by coherent simulations. In this work, we choose an iSWAP gate length balancing coherent error from the c-phase and incoherent error, arriving at a longer gate of order , roughly twice the CZ length. This additional c-phase error channel can be reduced by tailoring the hardware to the iSWAP, for example by increasing the anharmonicity or using a different gate architecture [41]. For this gate length, we measure an average c-phase of , yielding a Pauli error contribution to the gate of as shown in Fig. 4d, for a median inferred iSWAP Pauli error of (see SM sec. E3), larger the CZ gates used in the hexagonal implementation. This c-phase error accounts for roughly 19% of the detection probability in the bulk, as shown in Fig. 4e. Although our present device was optimized to realize CZ gates, the iSWAP c-phase error could be suppressed by increasing the anharmonicity in future designs.
Despite the lower gate fidelity on current hardware, one appeal of using iSWAP instead of CZ is the leakage generation and transport. As shown in Fig. 4f, the iSWAP’s on-resonance interaction is within the computational subspace and as a result does not directly populate non-computational states, minimizing leakage generation. When a higher excited state such as is present, it will transport to other qubits through the resonant interaction, and this transport can aid in leakage removal by spreading leakage from data qubits to measure qubits where a multi-level reset gate is performed each cycle. This behavior is in contrast to the CZ gate, which implements an on-resonance interaction with a precise timing calibrated to remove population from the state. Small errors in CZ calibration can thus generate leakage outside the computational space.
We probe the effect of leakage during memory experiments in Fig. 4h by measuring the population of excited states following the injection of a state on the center-most bulk data qubit after 10 cycles as shown in Fig. 4g, here using the distance-5 code. We repeat this experiment for three surface code implementations: the standard CZ implementation with and without DQLR, and the iSWAP implementation (which does not use DQLR). In a reference experiment where no leakage is injected (dotted lines), the standard surface code without DQLR reaches a steady state with oscillating average leakage between and per qubit due to the large leakage generation of the CZ gates. With DQLR, the average leaked population is considerably reduced as expected, oscillting between and . With the iSWAP implementation, the average leakage population reaches an oscillating steady state between and , limiting leakage well below the standard implementation without DQLR, indicating the superior leakage properties of the iSWAP gate compared with the CZ gate. Of all three, the standard CZ code with DQLR leads to the best leakage performance.
When the state is injected after a reset moment, all three implementations spike in their leaked population. The standard implementation without DQLR eliminates the extra leaked population within 10 cycles, whereas the DQLR version eliminate the leaked state within 2 cycles and the iSWAP in 4 cycles. This demonstrates that the iSWAP implementation can mitigate leakage well without explicit leakage removal, and adding DQLR to the iSWAP implementation could further improve its leakage characteristics.
Scaling and outlook
To compare surface code implementations on current hardware, in Fig. 5a we show contributions of each error component to the inverse scaling factor, , which can be expanded as a linear sum in physical error rates at first order (see SM sec. B). The total cycle times of the implementations are similar (hex: , walking: , iSWAP: ), allowing for direct comparison. These simulations go beyond the Pauli models used to construct the detector error budgets by including effects such as leakage and crosstalk. For all codes, these leakage and crosstalk contributions are still small relative to the two qubit gates, dynamical decoupling (DD), and readout error. The cphase contribution for the iSWAP code is significant, accounting for of the budget, and causes the iSWAP code’s simulated to be larger than the other three implementations.
For the hexagonal implementation, while the performance we measure already matches state of the art, tailoring the device to a honeycomb or brickwall lattice will reduce the number of connections needed from 2 to 1.5 per qubit (taking into account that qubits share connections). In principle, it will also reduce the number of potential error channels by reducing the complexity of the device. From a frequency optimization perspective, the lower connectivity reduces the number of variables, simplifying the optimization step. In Fig. 5b, we simulate the potential gain of tailoring the hardware to a hexagonal lattice by accounting for realistic device inhomogeneities in hardware parameters as well as TLS density with a frequency optimization step [27] (see SM sec. G). From these simulations we predict that simplifying the hardware by removing one coupler per qubit can lead to a improvement of the scaling factor . Such an increase in compounds favorably with a reduction in number of couplers needed, simplifying the design and number of control lines.
With these data, we have demonstrated state-of-the-art error suppression from distance-3 to distance-5 for the hexagonal, walking, and iSWAP surface code implementations. This proves that dynamic circuits are a viable approach to fault-tolerance, while also unlocking the ability to tailor the error correction circuit to hardware, presenting new opportunities for coupler design, multiplexing, and leakage removal intrinsic to the operation of QEC itself. In addition, our budgeting and leakage results provide direct insight into how using dynamic circuits uncovers new tradeoffs for realizing quantum error correction. We envision that other unexplored avenues for dynamic circuits may present further benefits, and our experiments solidify the prospects for the co-design of QEC and hardware as the state of the art for quantum error correction at scale.
Contributions
A. Morvan, M. McEwen, and A. Eickbusch conceived and led the project. A. Morvan and A. Eickbusch performed the calibrations and measurements. A. Morvan, A. Eickbusch and M. McEwen wrote the manuscript with input from all the authors. A. Morvan, A. Eickbusch, V. Sivak, M. McEwen, J. Atalaya and J. Claes wrote the supplements. M. McEwen, A. Eickbusch, A. Morvan, D. Bacon and C. Gidney designed the error correction circuits for the experiments. V. Sivak performed the reinforced learning decoding on all datasets and performed the simulation with gate frequency optimization and scaling to larger distance. A. Bourassa developed coherent error corrections for dynamic circuits. J. Atalaya, J. Claes, D. Kafri, A. Eickbusch, and M. McEwen performed the Pauli and hardware-accurate simulations. A. Eickbusch and A. Morvan developed the error budgeting for detectors. Z. Chen, A. Bengtsson, and A. Green performed initial bringup and calibration of the device. P. Klimov, W. Livingston, A. Pizzuto, and V. Sivak developed the algorithm specific frequency optimizations for these experiments. A. Bourassa, G. Roberts, K. Satzinger, and M. Neeley developed the infrastructure to translate QEC circuits to the hardware. A. Opremcak and A. Bengtsson performed calibration of the readout and dynamical decoupling. K. Miao and N. Zobrist performed the calibration of the reset and Leakage removal operations. C. Warren, and J. Gross developed the c-phase characterization of the iSWAP gate. All authors contributed to building the hardware and software, and writing the manuscript.
Acknowledgements
We thank Michel Devoret for extensive discussion on the manuscript.
Corresponding authors
Correspondence and requests for materials should be addressed to Alexis Morvan (amorvan@google.com), Alec Eickbusch (aleceickbusch@google.com) and Matt McEwen (mmcewen@google.com).
Competing Interests
The authors declare no competing interests.
Code and data availability
The datasets generated and analysed for this study are available at https://doi.org/10.5281/zenodo.14238907.
Google Quantum AI and Collaborators
Alec Eickbusch, Matt McEwen, Volodymyr Sivak1, Alexandre Bourassa1, Juan Atalaya1, Jahan Claes1, Dvir Kafri1, Craig Gidney1, Christopher W. Warren1, Jonathan Gross1, Alex Opremcak1, Nicholas Zobrist1, Kevin C. Miao1, Gabrielle Roberts1, Kevin J. Satzinger1, Andreas Bengtsson1, Matthew Neeley1, William P. Livingston1, Alex Greene1, Rajeev Acharya1, Laleh Aghababaie Beni1, Georg Aigeldinger1, Ross Alcaraz1, Trond I. Andersen1, Markus Ansmann1, Frank Arute1, Kunal Arya1, Abraham Asfaw1, Ryan Babbush1, Brian Ballard1, Joseph C. Bardin1, 3, Alexander Bilmes1, Jenna Bovaird1, Dylan Bowers1, Leon Brill1, Michael Broughton1, David A. Browne1, Brett Buchea1, Bob B. Buckley1, Tim Burger1, Brian Burkett1, Nicholas Bushnell1, Anthony Cabrera1, Juan Campero1, Hung-Shen Chang1, Ben Chiaro1, Liang-Ying Chih1, Agnetta Y. Cleland1, Josh Cogan1, Roberto Collins1, Paul Conner1, William Courtney1, Alexander L. Crook1, Ben Curtin1, Sayan Das1, Alexander Del Toro Barba1, Sean Demura1, Laura De Lorenzo1, Agustin Di Paolo1, Paul Donohoe1, Ilya K. Drozdov1, 2, Andrew Dunsworth1, Aviv Moshe Elbag1, Mahmoud Elzouka1, Catherine Erickson1, Vinicius S. Ferreira1, Leslie Flores Burgos1, Ebrahim Forati1, Austin G. Fowler1, Brooks Foxen1, Suhas Ganjam1, Gonzalo Garcia1, Robert Gasca1, Élie Genois1, William Giang1, Dar Gilboa1, Raja Gosula1, Alejandro Grajales Dau1, Dietrich Graumann1, Tan Ha1, Steve Habegger1, Michael C. Hamilton1, 4, Monica Hansen1, Matthew P. Harrigan1, Sean D. Harrington1, Stephen Heslin1, Paula Heu1, Oscar Higgott1, Reno Hiltermann1, Jeremy Hilton1, Hsin-Yuan Huang1, Ashley Huff1, William J. Huggins1, Evan Jeffrey1, Zhang Jiang1, Xiaoxuan Jin1, Cody Jones1, Chaitali Joshi1, Pavol Juhas1, Andreas Kabel1, Hui Kang1, Amir H. Karamlou1, Kostyantyn Kechedzhi1, Trupti Khaire1, Tanuj Khattar1, Mostafa Khezri1, Seon Kim1, Bryce Kobrin1, Alexander N. Korotkov1, Fedor Kostritsa1, John Mark Kreikebaum1, Vladislav D. Kurilovich1, David Landhuis1, Tiano Lange-Dei1, Brandon W. Langley1, Kim-Ming Lau1, Justin Ledford1, Kenny Lee1, Brian J. Lester1, Loïck Le Guevel1, Wing Yan Li1, Alexander T. Lill1, Aditya Locharla1, Erik Lucero1, Daniel Lundahl1, Aaron Lunt1, Sid Madhuk1, Ashley Maloney1, Salvatore Mandrà1, Leigh S. Martin1, Orion Martin1, Cameron Maxfield1, Jarrod R. McClean1, Seneca Meeks1, Anthony Megrant1, Reza Molavi1, Sebastian Molina1, Shirin Montazeri1, Ramis Movassagh1, Michael Newman1, Anthony Nguyen1, Murray Nguyen1, Chia-Hung Ni1, Logan Oas1, Raymond Orosco1, Kristoffer Ottosson1, Alex Pizzuto1, Rebecca Potter1, Orion Pritchard1, Chris Quintana1, Ganesh Ramachandran1, Matthew J. Reagor1, David M. Rhodes1, Eliott Rosenberg1, Elizabeth Rossi1, Kannan Sankaragomathi1, Henry F. Schurkus1, Michael J. Shearn1, Aaron Shorter1, Noah Shutty1, Vladimir Shvarts1, Spencer Small1, W. Clarke Smith1, Sofia Springer1, George Sterling1, Jordan Suchard1, Aaron Szasz1, Alex Sztein1, Douglas Thor1, Eifu Tomita1, Alfredo Torres1, M. Mert Torunbalci1, Abeer Vaishnav1, Justin Vargas1, Sergey Vdovichev1, Guifre Vidal1, Catherine Vollgraff Heidweiller1, Steven Waltman1, Jonathan Waltz1, Shannon X. Wang1, Brayden Ware1, Travis Weidel1, Theodore White1, Kristi Wong1, Bryan W. K. Woo1, Maddy Woodson1, Cheng Xing1, Z. Jamie Yao1, Ping Yeh1, Bicheng Ying1, Juhwan Yoo1, Noureldin Yosri1, Grayson Young1, Adam Zalcman1, Yaxing Zhang1, Ningfeng Zhu1, Sergio Boixo1, Julian Kelly1, Vadim Smelyanskiy1, Hartmut Neven1, Dave Bacon1, Zijun Chen1, Paul V. Klimov1, Pedram Roushan1, Charles Neill1, Yu Chen1, Alexis Morvan.
Google Quantum AI, Santa Barbara, CA 93117, USA
Department of Physics, University of Connecticut, Storrs, CT, USA
Department of Electrical and Computer Engineering, University of Massachusetts, Amherst, MA
Department of Electrical and Computer Engineering, Auburn University, Auburn, AL, USA
‡ These authors contributed equally to this work.